# Bounding causal effects with an unknown mixture of informative and non-informative missingness

### v3

@@ 1 Introduction
Causal inference requires that the intervention precede the outcome, imposing a necessary temporal ordering. However, in practice, some outcomes may be unobservable because units drop out of the study or become inaccessible to the researcher for other reasons during the period between the intervention and the outcome measurement. This issue can arise in virtually any type of study, even in randomized control trials. When outcomes are missing, identification assumptions are necessary to relate causal estimands to observable data quantities. A commonly adopted assumption is the “missing at random” (MAR) mechanism, under which outcome missingness is independent of the value of the outcome itself, conditional on all other observed data. Thus, any association between missingness and the outcome is fully explained by observable characteristics, enabling point identification of causal effects (see, e.g., williamson2012doubly; chaudhuri2016gmm; metten2022inverse). If instead missingness is a function of the outcomes themselves, or unobserved variables that determine both missingness and the outcomes, then the outcomes are said to be “missing not at random” (MNAR), and commonly targeted causal effects, such as the average treatment effect, are not in general identifiable.
@@
Unfortunately, testing whether MAR holds versus MNAR is not possible, and MAR is often implausible in practice. For example, in randomized trials examining disease progression or mortality, unobserved deterioration in health could directly lead to both study dropout and an adverse outcome. These issues are even more pronounced in observational studies utilizing electronic health records or insurance claims data, which typically feature higher rates of missingness. In this setting, outcomes such as diagnoses or procedures might not be documented if patients seek care elsewhere or forgo treatment altogether. Additionally, the reasons for missingness in such settings are often unknown yet plausibly linked to critical downstream outcomes, such as loss of insurance coverage. Furthermore, when missingness constitutes a competing event, an outcome that prevents the primary outcome from occurring, MAR is violated, and traditional causal estimands such as the average treatment effect may not be interpretable (stensrud2022separable).
@@
In contrast to the substantial literature addressing potential violations of the “ignorability” assumption, which states that treatment assignment is independent of potential outcomes conditional on observed covariates, there has been relatively limited exploration of MAR violations in causal inference. While ignorability holds by design in a randomized controlled trial, it generally does not hold in an observational study. As a result, a substantial body of literature has proposed sensitivity analyses that do not generally point identify causal effects, but instead provide bounds in cases where ignorability may not hold (liu2013introduction). However, less work has extended these methods to address violations of MAR when outcomes are missing.
@@
This paper addresses this gap by extending some of these methods, particularly those proposed by luedtke2015statistics, to accommodate scenarios in which MAR may be violated but ignorability still holds. We propose a novel sensitivity framework where missingness consists of an unknown mixture of informative (non-ignorable) and non-informative (ignorable) components. Our proposed framework yields general bounds for the average treatment effect under MAR violations. To enhance practical utility, we derive narrower bounds by specifying sensitivity parameters – such as the proportion of informative missingness – that link observable and unobservable data quantities. All bounds may be inverted to determine “tipping point” values of sensitivity parameters at which observed associations vanish, which may be preferred in situations where sensitivity parameters are difficult to specify a prior. We propose doubly-robust estimators (bang2005doubly) that incorporate flexible non-parametric and machine learning techniques (kennedy2022semiparametric; chernozhukov2018double) to estimate these bounds. Lastly, beyond bounds on the ATE, we provide bounds for alternative estimands more suitable in scenarios involving competing events, such as mortality, where violations of MAR are particularly pronounced. One such estimand that remains well-defined in the presence of competing events is the separable direct effect (stensrud2022separable).
@@ 1.1 Related work
The challenge of drawing valid inferences from data in the presence of missing outcomes has been extensively studied across various statistical settings. Missing outcomes routinely occur in longitudinal and time-to-event settings due to loss to follow up and other missingness. For example, survival analyses often make MAR assumptions (usually referred to as assuming non-informative censoring, or CAR) to point identify various functions of survival distributions (see, e.g., nicolaie2015vertical; do2017analysis; park2022semiparametric), and sensitivity analyses that relax the MAR assumption have been proposed (lipkovich2016sensitivity). However, a limitation of much of this literature is the absence of explicit causal interpretations of commonly estimated parameters; for instance, young2020causal demonstrated that hazard ratios generally lack clear causal interpretations. Importantly, our work does not directly address time-to-event outcomes but instead focuses on settings with outcomes measured at discrete time points or single measurements. Nevertheless, the problem of how to handle missing data, particularly when the missingness is due to a type of outcome that precludes the primary outcome from being measured (in the survival setting, a competing event), is directly relevant to this work.
@@
While a subset of the survival analysis literature explicitly addresses causal effects (see, e.g., robins1986new; frangakis2002principal; young2020causal), the primary focus has been on point identification and estimation young2020causal; stensrud2022separable. In contrast, our focus is specifically on deriving non-parametric bounds and conducting sensitivity analyses to address uncertainties due to informative missingness. Existing approaches proposing bounds and sensitivity analyses have largely relied on parametric models(andrea2001methods). Additionally, although distinctions between administrative censoring (typically assumed MAR) and competing risks (potentially informative censoring events) are recognized in the literature (mitra2020analysis; zhou2023case; han2021multiple; do2017analysis; moreno2013regression; lau2009competing; nicolaie2015vertical; lau2011parametric; varadhan2010evaluating), they generally assume these missingness mechanisms can be explicitly distinguished. In our setting, the researcher does not know which missing data are informative, complicating the analysis.
@@
Some prior work has considered similar problems at a single time point. Similar to the survival literature, missing-at-random assumptions are common in this work to point identify causal effects (see, e.g., chang2023covariate; zhao2024adjust). By contrast, mattei2014identification define causal effects within principal strata and derive bounds for these effects under non-ignorable missingness, though their approach relies on the availability of a binary instrumental variable separate from treatment assignment. jiang2022multiply introduce multiply-robust estimators based on a “principal ignorability” assumption, enabling flexible machine learning methods for estimation; however, this assumption is strong and may not hold in many practical scenarios. scharfstein2003incorporating propose Bayesian sensitivity analyses that summarize uncertainty in causal effects from non-ignorable missing outcomes but require explicit parametric assumptions about the missingness mechanism.
@@
Other researchers have jointly modeled missing outcomes and non-compliance in randomized experiments, leveraging randomization as an instrumental variable (frangakis1999addressing; baker2000; o2005likelihood; zhouandli2006). Non-parametric bounding techniques have also been developed for average treatment effects with binary outcomes in randomized contexts (horowitz2000nonparametric; taylor2006; gabriel2023nonparametric). We extend this non-parametric bounding approach to a broader context encompassing both randomized and observational studies. Our proposed bounds generalize previously derived results (such as those presented in horowitz2000nonparametric and equation (1) of gabriel2023nonparametric). We illustrate that without additional assumptions or constraints, these general bounds can often be overly wide, limiting their practical utility in scenarios with substantial missingness. Consequently, we provide strategies for incorporating subject matter expertise to yield more informative bounds.
@@
Finally, our approach shares some broad conceptual similarities with the sensitivity analysis framework of bonvini2022sensitivity. This paper proposes a sensitivity analysis for no unmeasured confounding assumption by positing that some set proportion of the observed data are confounded. Specifically, they define a latent indicator $S$ such that when $S=1$, no unmeasured confounding holds, and when $S=0$ it does not. However, our proposed approach is not simply an extension of this framework. For example, to extend bonvini2022sensitivity to our setting, we would instead define a latent indicator $S$ that defines a subset of units where MAR holds and a corresponding subset where it does not. However, our sensitivity analysis instead makes assumptions about the proportion of informatively missing observations among the missing observations – not the proportion of units where MAR holds. To be clear, even for observations without informatively missing outcomes, MAR does not necessarily hold. Extending the method from bonvini2022sensitivity to our setting ultimately requires substantively different starting assumptions, although extending this method directly would be an interesting avenue for future research.
@@ 2.1 Data, notation, and definitions
We assume that missingness is a mixture of two mutually exclusive types [^2] Technically, our results can be shown to hold without mutual exclusivity. However, without this requirement, the case where $U_{I}=1$ and $U_{NI}=1$ is difficult to conceptualize – either the missingness was informative or it was not. Therefore, we invoke mutual exclusivity for conceptual clarity., defined by unobserved indicators $U_{NI}$ and $U_{I}$. The observed missingness indicator $C$ is a composite of these variables given by $C=1-(1-U_{NI})(1-U_{I})=\mathbb{I}\{U_{NI}=1\text{ or }U_{I}=1\}$. $U_{NI}\in\{0,1\}$ represents non-informative or administrative missingness, while $U_{I}\in\{0,1\}$ represents informative missingness, which can be thought of as an intermediate outcome or a mediator of the effect of $A$ on $Y$. Missingness by $U_{NI}$ is non-informative when assumption (1) [^3] We thank an anonymous reviewer for observing that our submission used a stronger non-informative missingness assumption than was necessary for our later results to hold. holds:
@@
For example, in a clinical trial evaluating the effect of an antipsychotic drug on symptoms of psychosis and safety events such as increases in glycosylated hemoglobin (A1c), which may herald development of type 2 diabetes (for simplicity, hereafter we just refer the outcome as diabetes risk), participants might relocate and lose contact with trial administrators. Such administrative missingness can reasonably be assumed to be independent of diabetes risk, conditional on observed covariates, and thus is non-informative. By contrast, we place no restrictions on the dependence between $Y$ and $U_{I}$ beyond their temporal ordering. This permits informative missingness, where $U_{I}$ may not be conditionally independent of $Y$. In other words, $Y\cancel{\perp\!\!\!\perp}U_{I}|\mathbf{X},A$. Consider again the antipsychotic trial example. In some cases, participants may leave the trial before their outcomes are recorded due to medication side effects. Moreover, these side effects may also affect an individual’s diabetes risk. In such cases, missingness is not conditionally independent of the outcome $Y$ (and is therefore informative of the outcome).
@@
We introduce the following notation: let $e(\mathbf{x})=\mathbb{P}(A=1|\mathbf{X}=\mathbf{x})$ denote the probability of treatment (or propensity score); $\pi_{a}(\mathbf{x})=\mathbb{P}(C=1|\mathbf{X}=\mathbf{x},A=a)$ the probability of missingness in treatment arm $A=a$; and $\pi^{*}_{a}(\mathbf{x})=\mathbb{P}(U_{I}=1|\mathbf{X}=\mathbf{x},A=a)$ the probability of informative missingness in treatment arm $A=a$. Additionally, let $\mu_{a}(\mathbf{x})=\mathbb{E}(Y\mid\mathbf{X}=\mathbf{x},A=a,C=0)$ denote the conditional mean function in treatment arm $A=a$ among individuals with observed outcomes and $\mu_{a}^{\star}(\mathbf{x})=\mathbb{E}(Y\mid\mathbf{X}=\mathbf{x},A=a,U_{I}=1)$ the corresponding mean function among individuals informatively censored . We assume without loss of generality that $0<\mu_{a}(\cdot),\mu_{a}^{*}(\cdot)\leq 1$. This may be ensured by rescaling the bounded outcome or may occur by default with a binary outcome. Throughout, for any function $f$ of data $\bf O$}, $\mathbb{P}_{n}f$ denotes the empirical average $\frac{1}{n}\sum_{i=1}^{n}f({\bf O}_{i})$, $\mathbb{V}({\bf O})=\mathbb{E}\{({\bf O}-\mathbb{E}({\bf O}))({\bf O}-\mathbb{E}({\bf O}))^{{\sf\scriptscriptstyle{T}}}\}$, $\mathbb{I}(\cdot)$ is the indicator function, and we let $\|f({\bf O})\|^{2}=\int f({\bf o})^{2}d\mathbb{P}({\bf o})$ denote the squared $L_{2}$ norm.
@@ 2.2 Assumptions
Assumption 5 states that, in all strata defined by $\mathbf{X}$, outcomes have a positive probability of being observed. This assumption ensures sufficient overlap between the missing and non-missing observations to identify the ATE. These assumptions allow us to equate the ATE to a functional that does not involve counterfactual quantities.
@@
The proof is given in Appendix 9. However, this quantity is not identifiable in the data: given observed data $\bf O$}, we do not observe $U_{NI}$, only the composite indicator $C$, making the quantities $\pi_{a}^{*}(\mathbf{X})$ and $\mu_{a}^{\star}(\mathbf{X})$ inestimable without further assumptions. Nevertheless, as we will next show, it is possible to bound $\Psi_{0}$ under no additional assumptions, resulting in the general bounds we present in Section 3.1. Moreover, in section 3.2, we show that we may narrow these bounds by invoking additional assumptions that relate the unobserved quantities $\mu_{a}^{\star}$ and $\pi_{a}^{\star}$ to the observed quantities $\mu_{a}$ and $\pi_{a}$, possibly as a function of sensitivity parameters. Moreover, all bounds we will present are sharp under the stated motivating assumptions. In section 3.3, we propose sensitivity analyses that find which sensitivity parameters could explain away the naive $\widetilde{\Psi}$}, which may be useful when it is difficult to pre-specify values of the sensitivity parameters.

### v4

@@ 1 Introduction
Causal inference requires that the intervention precede the outcome, imposing a necessary temporal ordering. However, in practice, some outcomes may be unobservable because units drop out of the study or become inaccessible to the researcher for other reasons during the period between the intervention and the outcome measurement. This issue can arise in virtually any type of study, even in randomized controlled trials. When outcomes are missing, identification assumptions are necessary to relate causal estimands to observable data quantities. A commonly adopted assumption is the “missing at random” (MAR) mechanism, under which outcome missingness is independent of the value of the outcome itself, conditional on all other observed data. Thus, any association between missingness and the outcome is fully explained by observable characteristics, enabling point identification of causal effects (see, e.g., Williamson et al. (2012); Chaudhuri and Guilkey (2016); Metten et al. (2022)). If instead missingness is a function of the outcomes themselves, or unobserved variables that determine both missingness and the outcomes, then the outcomes are said to be “missing not at random” (MNAR), and commonly targeted causal effects, such as the average treatment effect, are not in general identifiable.
@@
Unfortunately, testing whether MAR holds versus MNAR is not possible, and MAR is often implausible in practice. For example, in randomized trials examining disease progression or mortality, unobserved deterioration in health could directly lead to both study dropout and an adverse outcome. These issues are even more pronounced in observational studies utilizing electronic health records or insurance claims data, which typically feature higher rates of missingness. In this setting, outcomes such as diagnoses or procedures might not be documented if patients seek care elsewhere or forgo treatment altogether. Additionally, the reasons for missingness in such settings are often unknown yet plausibly linked to critical downstream outcomes, such as loss of insurance coverage. Furthermore, when missingness constitutes a competing event, an outcome that prevents the primary outcome from occurring, MAR is violated, and traditional causal estimands such as the average treatment effect may not be interpretable Stensrud et al. (2022).
@@
In contrast to the substantial literature addressing potential violations of the “ignorability” assumption, which states that treatment assignment is independent of potential outcomes conditional on observed covariates, there has been relatively limited exploration of MAR violations in causal inference. While ignorability holds by design in a randomized controlled trial, it generally does not hold in an observational study. As a result, a substantial body of literature has proposed sensitivity analyses that do not generally point identify causal effects, but instead provide bounds in cases where ignorability may not hold Liu et al. (2013). However, less work has extended these methods to address violations of MAR when outcomes are missing.
@@
This paper addresses this gap by extending some of these methods, particularly those proposed by Luedtke et al. (2015), to accommodate scenarios in which MAR may be violated but ignorability still holds. We propose a novel sensitivity framework where missingness consists of an unknown mixture of informative (non-ignorable) and non-informative (ignorable) components. Our proposed framework yields general bounds for the average treatment effect under MAR violations. To enhance practical utility, we derive narrower bounds by specifying sensitivity parameters – such as the proportion of informative missingness – that link observable and unobservable data quantities. All bounds may be inverted or evaluated across a range of sensitivity parameters to determine “tipping point” values of sensitivity parameters at which observed associations vanish, which may be preferred in situations where sensitivity parameters are difficult to specify a priori. We propose doubly-robust estimators Bang and Robins (2005) that incorporate flexible non-parametric and machine learning techniques Kennedy (2022); Chernozhukov et al. (2018) to estimate these bounds. Lastly, beyond bounds on the ATE, we provide bounds for alternative estimands more suitable in scenarios involving competing events, such as mortality, where violations of MAR are particularly pronounced. One such estimand that remains well-defined in the presence of competing events is the separable direct effect Stensrud et al. (2022). We implement the proposed methods in the marbounds package for R, available at https://github.com/mrubinst757/marbounds.
@@ 1.1 Related work
The challenge of drawing valid inferences from data in the presence of missing outcomes has been extensively studied across various statistical settings. Missing outcomes routinely occur in longitudinal and time-to-event settings due to loss to follow up and other missingness. For example, survival analyses often make MAR assumptions (usually referred to as assuming non-informative censoring, or CAR) to point identify various functions of survival distributions (see, e.g., Nicolaie et al. (2015); Do and Kim (2017); Park et al. (2022)), and sensitivity analyses that relax the MAR assumption have been proposed Lipkovich et al. (2016). However, a limitation of much of this literature is the absence of explicit causal interpretations of commonly estimated parameters; for instance, Young et al. (2020) demonstrated that hazard ratios generally lack clear causal interpretations. Importantly, our work does not directly address time-to-event outcomes but instead focuses on settings with outcomes measured at discrete time points or single measurements. Nevertheless, the problem of how to handle missing data, particularly when the missingness is due to a type of outcome that precludes the primary outcome from being measured (in the survival setting, a competing event), is directly relevant to this work.
@@
While a subset of the survival analysis literature explicitly addresses causal effects (see, e.g., Robins (1986); Frangakis and Rubin (2002); Young et al. (2020)), the primary focus has been on point identification and estimation Young et al. (2020); Stensrud et al. (2022). In contrast, our focus is specifically on deriving non-parametric bounds and conducting sensitivity analyses to address uncertainties due to informative missingness. Existing approaches proposing bounds and sensitivity analyses have largely relied on parametric models Andrea et al. (2001). Additionally, although distinctions between administrative censoring (typically assumed MAR) and competing risks (potentially informative censoring events) are recognized in the literature (see, e.g., Mitra et al. (2020); Zhou et al. (2023); Han et al. (2021); Do and Kim (2017); Moreno-Betancur and Latouche (2013); Lau et al. (2009); Nicolaie et al. (2015); Lau et al. (2011); Varadhan et al. (2010)), they generally assume these missingness mechanisms can be explicitly distinguished. In our setting, the researcher does not know which missing data are informative, complicating the analysis.
@@
Some prior work has considered similar problems at a single time point. Similar to the survival literature, missing-at-random assumptions are common in this work to point identify causal effects (see, e.g., Chang et al. (2023); Zhao and Ding (2024)). By contrast, Mattei et al. (2014) define causal effects within principal strata and derive bounds for these effects under non-ignorable missingness, though their approach relies on the availability of a binary instrumental variable separate from treatment assignment. Jiang et al. (2022) introduce multiply-robust estimators based on a “principal ignorability” assumption, enabling flexible machine learning methods for estimation; however, this assumption is strong and may not hold in many practical scenarios. Scharfstein et al. (2003) propose Bayesian sensitivity analyses that summarize uncertainty in causal effects from non-ignorable missing outcomes but require explicit parametric assumptions about the missingness mechanism.
@@
Other researchers have jointly modeled missing outcomes and non-compliance in randomized experiments, leveraging randomization as an instrumental variable (Frangakis and Rubin (1999); Baker (2000); O’Malley and Normand (2005); Zhou and Li (2006)). Non-parametric bounding techniques have also been developed for average treatment effects with binary outcomes in randomized contexts (Horowitz and Manski (2000); Taylor and Zhou (2006); Gabriel et al. (2023)). We extend this non-parametric bounding approach to a broader context encompassing both randomized and observational studies. Our proposed bounds generalize previously derived results, including those in Horowitz and Manski (2000) and equation (1) of Gabriel et al. (2023). We illustrate that without additional assumptions or constraints, these general bounds can often be overly wide, limiting their practical utility in scenarios with substantial missingness. Consequently, we provide strategies for incorporating subject matter expertise to yield more informative bounds.
@@
Finally, our approach shares some broad conceptual similarities with the sensitivity analysis framework of Bonvini and Kennedy (2022). This paper proposes a sensitivity analysis for no unmeasured confounding assumption by positing that some set proportion of the observed data are confounded. Specifically, they define a latent indicator $S$ such that when $S=1$, no unmeasured confounding holds, and when $S=0$ it does not. However, our proposed approach is not simply an extension of this framework. For example, to extend Bonvini and Kennedy (2022) to our setting, we would instead define a latent indicator $S$ that defines a subset of units where MAR holds and a corresponding subset where it does not. However, our sensitivity analysis instead makes assumptions about the proportion of informatively missing observations among the missing observations – not the proportion of units where MAR holds. To be clear, even for observations without informatively missing outcomes, MAR does not necessarily hold. Extending the method from Bonvini and Kennedy (2022) to our setting ultimately requires substantively different starting assumptions, although doing so would be an interesting avenue for future research.
@@ 2.1 Data, notation, and definitions
We assume that missingness is a mixture of two mutually exclusive types [^2] Technically, our results can be shown to hold without mutual exclusivity. However, without this requirement, the case where $U_{I}=1$ and $U_{NI}=1$ is difficult to conceptualize – either the missingness was informative or it was not. Therefore, we invoke mutual exclusivity for conceptual clarity., defined by unobserved indicators $U_{NI}$ and $U_{I}$. The observed missingness indicator $C$ is a composite of these variables given by $C=1-(1-U_{NI})(1-U_{I})=\mathbb{I}\{U_{NI}=1\text{ or }U_{I}=1\}$. $U_{NI}\in\{0,1\}$ represents non-informative or administrative missingness, while $U_{I}\in\{0,1\}$ represents informative missingness, which can be thought of as an intermediate outcome or a mediator of the effect of $A$ on $Y$. Missingness by $U_{NI}$ is non-informative when assumption (1) [^3] We thank an anonymous reviewer for observing that our initial submission used a stronger non-informative missingness assumption than was necessary for our later results to hold. holds:
@@
For example, in a clinical trial evaluating the effect of an antipsychotic drug on symptoms of psychosis and safety events such as increases in glycosylated hemoglobin (A1c), which may herald development of type 2 diabetes (for simplicity, hereafter we refer to the outcome as diabetes risk), participants might relocate and lose contact with trial administrators. Such administrative missingness can reasonably be assumed to be independent of diabetes risk, conditional on observed covariates, and thus is non-informative. By contrast, we place no restrictions on the dependence between $Y$ and $U_{I}$ beyond their temporal ordering. This permits informative missingness, where $U_{I}$ may not be conditionally independent of $Y$. In other words, $Y\cancel{\perp\!\!\!\perp}U_{I}|\mathbf{X},A$. Consider again the antipsychotic trial example. In some cases, participants may leave the trial before their outcomes are recorded due to medication side effects. Moreover, these side effects may also affect an individual’s diabetes risk. In such cases, missingness is not conditionally independent of the outcome $Y$ (and is therefore informative of the outcome).
@@
We introduce the following notation: let $e(\mathbf{x})=\mathbb{P}(A=1|\mathbf{X}=\mathbf{x})$ denote the probability of treatment (or propensity score), where we at times use $e_{1}(\mathbf{X})$ to denote $e(\mathbf{X})$ and $e_{0}(\mathbf{X})$ to denote $1-e(\mathbf{X})$; $\pi_{a}(\mathbf{x})=\mathbb{P}(C=1|\mathbf{X}=\mathbf{x},A=a)$ the probability of missingness in treatment arm $A=a$; and $\pi^{*}_{a}(\mathbf{x})=\mathbb{P}(U_{I}=1|\mathbf{X}=\mathbf{x},A=a)$ the probability of informative missingness in treatment arm $A=a$. Additionally, let $\mu_{a}(\mathbf{x})=\mathbb{E}(Y\mid\mathbf{X}=\mathbf{x},A=a,C=0)$ denote the conditional mean function in treatment arm $A=a$ among individuals with observed outcomes and $\mu_{a}^{\star}(\mathbf{x})=\mathbb{E}(Y\mid\mathbf{X}=\mathbf{x},A=a,U_{I}=1)$ the corresponding mean function among individuals informatively censored. We assume without loss of generality that $0<\mu_{a}(\cdot),\mu_{a}^{\star}(\cdot)\leq 1$. This may be ensured by rescaling the bounded outcome or may occur by default with a binary outcome. Throughout, for any function $f$ of data $\bf O$}, $\mathbb{P}_{n}f$ denotes the empirical average $\frac{1}{n}\sum_{i=1}^{n}f({\bf O}_{i})$, $\mathbb{V}({\bf O})=\mathbb{E}\{({\bf O}-\mathbb{E}({\bf O}))({\bf O}-\mathbb{E}({\bf O}))^{{\sf\scriptscriptstyle{T}}}\}$, $\mathbb{I}(\cdot)$ is the indicator function, and we let $\|f({\bf O})\|^{2}=\int f({\bf o})^{2}d\mathbb{P}({\bf o})$ denote the squared $L_{2}(\mathbb{P})$ norm.
@@ 2.2 Assumptions
Assumption 5 states that, in all strata defined by $\mathbf{X}$, outcomes have a positive probability of being observed. This assumption ensures sufficient overlap between the missing and non-missing observations to identify the observed data functionals that appear below. These assumptions allow us to equate the ATE to a functional that does not involve counterfactual quantities.
@@
The proof is given in Appendix 10. However, this quantity is not identifiable in the data: given observed data $\bf O$}, we do not observe $U_{NI}$, only the composite indicator $C$, making the quantities $\pi_{a}^{\star}(\mathbf{X})$ and $\mu_{a}^{\star}(\mathbf{X})$ inestimable without further assumptions. Nevertheless, as we will next show, it is possible to bound $\Psi_{0}$ under no additional assumptions, resulting in the general bounds we present in Section 3.1. Moreover, in section 3.2, we show that we may narrow these bounds by invoking additional assumptions that relate the unobserved quantities $\mu_{a}^{\star}$ and $\pi_{a}^{\star}$ to the observed quantities $\mu_{a}$ and $\pi_{a}$, possibly as a function of sensitivity parameters. Moreover, all bounds we will present are sharp under the stated motivating assumptions. In section 3.4, we propose sensitivity analyses that find which sensitivity parameters could explain away the naive $\widetilde{\Psi}$}, which may be useful when it is difficult to pre-specify values of the sensitivity parameters.